How fast is a 10 kg box going to be moving from rest after 3 seconds if a 200 N force is applied to the left and a 50 N force is applied to the right?


1
Expert's answer
2021-05-05T17:53:23-0400

mass(m)=10kg

time(t)=3sec

F1=200N towards left

F2=50N towards right

since F1>F2So body will move towards left.

Applying newton's law of motion

F1>F2= ma----------------------------(1)

200-50= 10 v/t acceleration(a)= v/t

150= (10/3)v

v= 450/10

v=45 m/s (ANSWER)